OpenStudy (anonymous):

what is the answer for x+2/9=5

OpenStudy (anonymous):

\[\frac{x+2}{9}=5\] or \[x+\frac{2}{9}=5\]?

OpenStudy (doc.brown):

If your teacher follows bedmas or pemdas \[x+\frac{2}{9} = 5\]\[x = 5 - \frac{2}{9}\]\[x = 5(\frac{45}{45}) - \frac{2}{9}(\frac{5}{5})\]\[x = \frac{225}{45} - \frac{10}{45}\]\[x = \frac{215}{45} = \frac{43}{9}\]\[x = 4\frac{7}{9}\]